OpenStudy (anonymous):

x + 4)(x + 2) = A. x2 + 6x + 6 B. x2 + 8x + 8 C. x2 + 8x + 6 D. x2 + 6x + 8

OpenStudy (psymon):

x*x = x*2 = 4*x = 4*2 = It's just using FOIL. Multiply those 4 individual problems and then add up the results : )

OpenStudy (nurali):

(x + 4)(x + 2) (x+4)(x+2)=x*x+x*2+4*x+4*2 (x+4)(x+2)=x^2+6x+8

OpenStudy (lenovo):

\[(x+4)(x+2)=x(x+2)+4(x+2)=x ^{2}+2x+4x+8=x ^{2}+6x+8\]

OpenStudy (lenovo):

So the answer is option D

OpenStudy (cggurumanjunath):

d